OpenStudy (lamechildidc):

Please help..I need to pass this. What is the sum of the geometric sequence 1, 3, 9, ... if there are 11 terms? 29,524 55,987 87,381 88,573 I added everything up and got D. But, I searched the question to be sure..and others said A. Which confuses me. I need confirmation, thank you!

OpenStudy (anonymous):

⇒S=a×1−rn1−r substitute r with 3(we're multiplying 3 everytime), n with 11.

OpenStudy (anonymous):

a is the first term aka 1.

OpenStudy (roxy.girl):

sum of a geometric series = a X ( 1- r^n)/(1 - r) here a = 1 r = 3 and n=11 sum = 1 X 1 - 3^11/(1-3) = 1- 3^11/-2 =1- 177147 = - 177146/-2 =?

OpenStudy (lamechildidc):

88,573

OpenStudy (lamechildidc):

So, i was correct?

OpenStudy (anonymous):

yes

OpenStudy (roxy.girl):

Yes that is the right answer ^.^ Not sure about how A got in there..

OpenStudy (lamechildidc):

Okay thank you. Can you guys help with another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

oh could i have medal pls

OpenStudy (lamechildidc):

I'll put it in another question and if you help me I'll give you the medal. One sec